JET SPECTROSCOPY OF THE ν_{4} BAND OF HYDROGEN FLUORIDE TETRAMER IN THE 714 cm^{−1}REGION

\underline{THOMAS A. BLAKE} and STEVEN W. SHARPE, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, 3020 Q Avenue, Richland, WA 99352.

Using our slit-jet, infrared diode laser spectrometer, we have observed a rotationally resolved parallel band centered at 714 cm^{−1}in a pulsed molecular beam formed by expanding a mixture of 4% HF in helium through a 200 μm x 12 cm slit with a backing pressure of ca. 1000 Torr. A set of excited and ground state spectroscopic constants were fit simultaneously to a data set of 204 transitions using a symmetric top, semi-rigid rotor Hamiltonian. The fit constants are ν_{0} = 714.7849(1) cm^{−1}, B′ = 0.129614(5) cm^{−1}, (C′ − B′) − (C′′ − B′′) = 0.001339(6) cm^{−1}, D′_{J} = 4.9(2) × 10^{−7} cm^{−1}, D′_{JK} = −1.1(2) × 10^{−6} cm^{−1}, D′_{K} − D′′_{K} = −2.81(8) × 10^{−6} cm^{−1}, H′_{JKK} = −2.0(2) × 10^{−8} cm^{−1}, H′_{KKK} − H′′_{KKK} = −4.68(6) × 10^{−8} cm^{−1}, B′′ = 0.132059(5) cm^{−1}, D′′_{J} = 5.1(2) × 10^{−7} cm^{−1}, D′′_{JK} = −1.3(2) × 10^{−6} cm^{−1}, H′′_{JKK} = −6(2) × 10^{−9} cm^{−1}, and the RMS deviation of the obs - calcs is 0.00045 cm^{−1}. We have also observed two other bands in the 700 to 775 cm^{−1}region: one band is a strongly perturbed perpendicular band centered at 752 cm^{−1}with B′′ = 0.133 cm^{−1}and the other band is a perturbed parallel band centered at 741 cm^{−1}with B′′ = 0.0755 cm^{−1}. For hydrogen fluoride tetramer and pentamer, theory predicts (Maerker, et al.) ^{a} planar, oblate symmetric top structures at their global minima with B_{e} values of 0.136 and 0.0767 cm^{−1}for (HF)_{4} (C_{4h} symmetry) and for (HF)_{5} (C_{5h} symmetry), respectively. We have tentatively assigned the 714 cm^{−1}band as the ν_{4} (A_{u}) HF out-of-plane torsion of the tetramer, the 752 cm^{−1}band as the ν_{13} (E_{u}) HF in-plane torsion of the tetramer, and the 741 cm^{−1}band as the ν_{4} (A′′) HF out-of-plane torsion of the pentamer.

aC. Maerker, P. R. Schlyer, K. R. Liedl, T. -K. Ha, M. Quack, M. A. Suhm, J. Comp. Chem. 18 1695 (1997).